An asteroid of mass 58,000 kg carrying a negative charge of 15.0 ?C is 180 m from a second asteroid of mass 52,000 kg carrying a negative charge of 11.0 ?C . What is the net force the asteroids exert upon each other? (G = 6.673
Answer #1

i have a same question with solution, just you need to change the figures according to your question.

An asteroid of mass 59,000 kg carrying a negative charge of 12.0 μC is 170 m from a second asteroid of mass 58,000 kg carrying a negative charge of 15.0 μC . What is net force the asteroids exert upon each other?

Solution:

gravitational force = G M m / d^2 =

= 6.67x10^-11 * 59000 * 58000 / 170^2 =

=7.8978 x 10^-6 Newtons

elec force = k Q q / d^2 =

= 8.99x10^9 * 0.000012 * 0.000015 / 170^2 =

= 55.9931 x 10^-6 Newtons

gravity is attractive, elec force is repulsive, so total force is repulsive and equal to

55.9931 - 7.8978 = 48.1 x 10^-6 Newtons
Answer #2
0.000040 N,
Key concepts to use: FE = (K|q1||q2|)/r2 and Fg = (Gm1m2)/r2